Suppose one basket has five white balls and seven black balls. A second basket has three...

  1. Suppose one basket has five white balls and seven black balls. A second basket has three white balls and twelve black balls. We will flip a fair coin and if it lands “heads” a ball from basket 1 is selected, and if it lands “tails” a ball from basket 2 is selected. Suppose that we have done this and that a white ball was selected, what is the probability that the coin landed tails?
Answer #1

Basket 1: 5 white, 7 black

Thus, P(white/basket1) = 5/12

Basket 2: 3 white, 12 black

Thus, P(white/basket2) = 3/15 = 1/5

We will flip a fair coin and if it lands “heads” a ball from basket 1 is selected, and if it lands “tails” a ball from basket 2 is selected

Thus, P(selecting basket 1) = P(selecting basket 2) = 1/2

P(tails/ white ball was selected )

= P(basket 2 / white ball was selected)

= P(basket 2 and white ball was selected) / P(white ball was selected)

= (1/5) / (1/5 + 5/12)

= (1/5) / (37/60)

= 12/37

= 0.3243